# Transverse vortices induced by modulated granular shear flows of elongated particles

### v1

@@ Abstract
We have performed DEM simulations of elongated grains in a shear cell for various particle aspect ratios and contact frictions, with an additional heterogeneous force perturbation in the flow direction. For a perturbation in the form of a single Fourier mode, we show that the response of the system consists of transverse secondary flows that average onto a pattern of four vortices. We have also theoretically studied this phenomenon by generalizing the granular rheology $\mu(I)$ to the case of elongated grains and computing the linear response to such a perturbation. Even if the agreement between theory and simulations remains qualitative only, we can reproduce and understand the inversion of the vortex pattern when the cell aspect ratio is increased from a vertically to a horizontally elongated cell shape, emphasizing the key role of the second normal stress difference as well as the cell geometry.
@@ I Introduction
Secondary flows in granular systems have been observed in a variety of configurations. Vortex-like convective rolls have been studied in gravity-driven chute flows, where regions with very different densities develop, especially when highly agitated gas-like domains appear with considerably decreased density Forterre and Pouliquen (2001); Börzsönyi et al. (2009); Brodu et al. (2013, 2015). Taylor vortex-like structures have also been seen in a granular flow Conway et al. (2004), which are analogous to the well-known instability of liquid flows in the Taylor-Couette geometry Taylor (1923). The formation of heaps on the top surface of the granular material was also observed in a similar geometry Cabrera and Polanía (2020). In slower flows, where density variations are much smaller, density-driven secondary convection was studied in a cylindrical Couette device and as well as numerically in a split bottom shear cell Krishnaraj and Nott (2016); Dsouza et al. (2017); Dsouza and Nott (2021). Recently, a slow transverse flow was created by a bulldozing mechanism in driving granular material with a conveyor belt Einav et al. (2024).
@@
Secondary flows are also present in flows of granular suspensions Guazzelli and Pouliquen (2018), which share part of their rheological behavior with dry granular systems Boyer et al. (2011); Trulsson et al. (2012). Secondary flows, possibly associated with migration of particles, are known to be induced by normal stress differences Ramachandran and Leighton (2008), as exemplified by the Weissenberg effect Yamamoto (1958); Oldroyd (1958); Kundu (1973); Elson et al. (1982) with viscoelastic fluids such as polymers. Normal stress differences in granular suspensions have been measured in experiments analyzing the curvature of the free surface of flows in a tilted trough Couturier et al. (2011), or recording stress profiles in a torsional flow between discs Dbouk et al. (2013), as well as in simulations Seto and Giusteri (2018). Interestingly, they get enhanced values for elongated particles such as fibers or dimers Snook et al. (2014); Bounoua et al. (2016); Butler and Snook (2018); Mari (2020). The free-surface curvature of channelized gravity-driven flows of dry grains has also been associated with the presence of normal stress differences McElwaine et al. (2012); Gadal et al. (2025), especially in the context of the formation of levees Félix and Thomas (2004); Deboeuf et al. (2006); Rocha et al. (2019).
@@
The flow of dense granular systems has been largely studied over the last two decades, in particular developing an empirical frictional rheology, known as the $\mu(I)$ rheology GdR MiDi (2004); da Cruz et al. (2005); Jop et al. (2006); Forterre and Pouliquen (2008); Jop (2015). Starting from the idealized case of systems of mono-sized hard spheres, this rheology has been extended in various ways, for example, accounting for grain cohesion Rognon et al. (2006); Mandal et al. (2020); Blatny et al. (2024), grain softness and contact friction Favier de Coulomb et al. (2017) or active (self-propelled) grains Peshkov et al. (2016). An important piece of work has also been done to properly accommodate, or regularize, this rheology in Navier-Stokes solvers Barker et al. (2015, 2017); Goddard and Lee (2018). In case of the presence of strong gradients, a non-local rheology has been developed Pouliquen and Forterre (2009); Bouzid et al. (2013); Henann and Kamrin (2013), for which an additional variable, the fluidity, has to be defined and modeled Bouzid et al. (2015); Zhang and Kamrin (2017); Poon et al. (2023). Finally, the kinetic theory for granular gases, pushed to the dense limit, recovers the features of the granular rheology and its non-local extensions Berzi (2024).
@@
For our present purpose, a key rheological aspect is to account for grain elongation, as e.g. with rods Börzsönyi et al. (2012a); Guo et al. (2013); Berzi et al. (2022), multisphere particles Börzsönyi et al. (2012b); Liu et al. (2024); Rahim et al. (2024), spherocylinders or ellipses Nagy et al. (2017); Trulsson (2018); Nagy et al. (2020). These studies have shown that, in a similar way to granular suspensions, these systems display enhanced normal stress differences. In our wish to understand the processes at work in the heaping problem, and to relate secondary transverse vortices to these normal stress differences, the idea is to investigate the flow of grains of arbitrary elongation aspect ratio $\alpha$ in the simpler configuration of the shear cell, where a non-homogeneous forcing is implemented in the flow direction. This forcing will be of arbitrarily small amplitude $\epsilon$ so that we can compute the response of the system at the linear order in this perturbation.
@@
The structure of the paper is the following: in Section II we propose a generalized granular rheology in the case of elongated grains and study the case of a plane Couette flow. In Sections III and IV we analytically compute the linear response of a shear flow to a nonhomogeneous forcing in the flow direction. We introduce the discrete element method (DEM) and numerical setup to simulate such granular flows in Section V. In Section VI we present and discuss the results of these simulations in comparison to the theoretical expectations.
@@ II.1 Reference isotropic case
Assuming pressure isotropy, it has been proposed by Jop et al. Jop et al. (2006) that the granular rheology, which relates the stress tensor $\sigma_{ij}$ to the velocity field $u_{i}$ takes the form:
@@
where $\Pi=\tfrac{1}{3}\sigma_{ii}$ is the pressure and $\dot{\gamma}_{ij}=\dot{\Gamma}_{ij}+\dot{\Gamma}_{ji}$ with $\dot{\Gamma}_{ij}=\partial_{i}u_{j}$ is the shear rate. The norm of the shear rate is here defined as $\dot{\gamma}^{2}=\tfrac{1}{2}\dot{\gamma}_{ij}\dot{\gamma}_{ij}$. The effective friction $\mu$ is a function of the inertial number $I=\dot{\gamma}d/\sqrt{\Pi/\rho}$, where $d$ is the grain size and $\rho$ the grain mass density, and is empirically determined by measurements on experimental or numerical data. Although this rheology also comes with a second constitutive relation describing the behavior of the packing fraction $\phi(I)$, this expression (1) must be consistent with the definition of $\Pi$, incompressibility: $\dot{\gamma}_{ii}=0$. We will eliminate this possible issue by assuming later the quasi-static limit $I\to 0$.
Figure 1: Schematic drawing of the studied system. (a) the simulated cell has a dimension $L_{x}\times L_{y}\times L_{z}$. It is confined by $\sigma_{yy}$ in the vertical direction. Shear is applied in the $x$-direction, plus (b) an additional external force $F_{x}(y,z)$ in the $x$-direction (its magnitude and sign are shown by the surface and the green arrows). This results in the flow pattern shown in (c), four vertices in the velocity field in the $y-z$ plane.
@@ II.2 Generalized constitutive relation
We would like to generalize this rheology (1) to the case of elongated grains, for which finite normal stress differences $N_{1}=\sigma_{xx}-\sigma_{yy}$ and $N_{2}=\sigma_{yy}-\sigma_{zz}$ arise. To do so, one must notice that the following symmetries are required:
  • The normal stress differences depend on the direction of the velocity gradient and the flow, so they must change with the transformation $\dot{\bm{\Gamma}}\rightarrow\dot{\bm{\Gamma}}^{T}$. This means they must depend explicitly on $\dot{\Gamma}_{ij}$, as $\dot{\gamma}_{ij}$ is symmetric.
  • The normal stress differences are invariant to the reversal of the flow direction ($\dot{\Gamma}_{ij}\rightarrow-\dot{\Gamma}_{ij}$), so they can only depend on even order terms of the shear rate.
  • The new term should have a trace of zero, to keep the definition of $\Pi$ correct.
The simplest relation that fulfills these requirements is:
@@
or, without index notation, in a coordinate-independent form:
@@
with $\dot{\gamma}^{2}=\tfrac{1}{2}\operatorname{Tr}(\dot{\bm{\gamma}}\dot{\bm{\gamma}}^{T})$. The last two terms must have the same factor $\eta$ to ensure that the stress tensor is symmetric. The coefficients $\mu$, $\nu_{1}$, $\nu_{2}$, and $\eta$ depend on the inertial number $I$, but also on microscopic parameters such as the grain aspect ratio $\alpha$ or the grain contact friction $\mu_{p}$ Nagy et al. (2017, 2020); Srivastava et al. (2021). This rheological form is similar to a Rivlin–Ericksen stress closure, as proposed for granular flows in McElwaine et al. (2012); Srivastava et al. (2021); Kim and Kamrin (2023); Gadal et al. (2025).
@@ II.3 Plane Couette flow
The granular rheology is usually studied and calibrated in the geometry of a plane Couette flow. In this case, we consider a cell sheared in the $x$-direction with the gradient in the $y$ (‘vertical’) direction, and $z$ being the neutral (‘transverse’) direction. This cell has dimensions $L_{x}$, $L_{y}$, and $L_{z}$ along those three axes, and we take the origin of the coordinates at the center of the box. The grains in this cell are subjected to a given pressure $\Pi_{0}$. The flow in the homogeneous case, which will be the base state of the calculation later on, is characterized by a linear velocity profile, associated with a reference shear rate $\dot{\gamma}_{0}$:
@@
In this configuration for which only $\dot{\Gamma}_{yx}\neq 0$, both $\dot{\bm{\Gamma}}_{0}^{2}$ and $(\dot{\bm{\Gamma}}_{0}^{T})^{2}$ are zero, so that the second line of (3) disappears and the coefficient $\eta$ cannot be measured in such a configuration. To measure it, a configuration with extensional or compressional flows is required. It will then be left adjustable in the following. In contrast, $\nu_{1}$ and $\nu_{2}$ are related to the calibrated first and second normal stress differences as:
@@
whose values are taken from Nagy et al. Nagy et al. (2020) (see also Fig. A1 in Appendix A). The traces are related as:
@@
As mentioned above, we will restrict to the quasi-static limit for which $I_{0}=\dot{\gamma}_{0}d/\sqrt{\Pi_{0}/\rho}\ll 1$. Under this assumption, the stress tensor in this base state finally reads:
@@
where $\pi_{c}=1-\tfrac{1}{3}(\nu_{1c}+\nu_{2c})$, and where the subscripts ‘c’ indicate that $\mu$, $\nu_{1}$ and $\nu_{2}$ are taken constants (independent of the shear rate, but still functions of $\alpha$ and $\mu_{p}$, see Appendix A).
@@ III Conservation equations and perturbation
The equations that govern the dynamics of the system are the conservation of mass and momentum. Assuming incompressibility ($\phi\sim\phi_{c}$ is constant), these equations read:
@@
Here $b_{i}$ is an external body force density. It is an arbitrary perturbation. We take it in the following form:
@@
where $\epsilon$ is its vanishing dimensionless amplitude and $f$ its normalized profile in the $y-z$ plane. We take no component along the $y$- and $z$-directions.
Assuming that $\epsilon\ll 1$, the response of velocity and pressure to this perturbation can be written in linear order with respect to the base state as:
@@
where $U_{i}$ and $P$ are unknown functions that we wish to relate to $f$. With these expressions, one can compute the shear rate:
@@
the symmetric shear rate:
@@
and the second-order terms:
@@
Even if $\dot{\bm{\Gamma}}^{2}$ and $(\dot{\bm{\Gamma}}^{T})^{2}$ now do not vanish, their traces are still zero to the second order in $\epsilon$, so that:
@@
Finally, the stress tensor is:
@@
where, using (3) and the above expressions, we obtain:

### v2

@@ Abstract
We perform discrete element method (DEM) simulations of elongated grains in a shear cell for various particle aspect ratios and contact frictions, with an additional heterogeneous force perturbation in the flow direction. For a perturbation in the form of a single Fourier mode, we show that the response of the system consists of transverse secondary flows that average onto a pattern of four vortices. We also theoretically studied this phenomenon by generalizing the granular rheology $\mu(I)$ to the case of elongated grains and computing the linear response to such a perturbation. Even if the agreement between theory and simulations remains qualitative only, we can reproduce and understand the inversion of the vortex pattern when the cell aspect ratio is increased from a vertically to a horizontally elongated cell shape, emphasizing the key role of the second normal stress difference as well as the cell geometry.
@@ I Introduction
Secondary flows in granular systems have been observed in a variety of configurations. Vortex-like convective rolls have been studied in gravity-driven chute flows, where regions with very different densities develop, especially when highly agitated gas-like domains appear with considerably decreased density Forterre and Pouliquen (2001); Börzsönyi et al. (2009); Brodu et al. (2013, 2015). Taylor vortex-like structures have also been seen in a granular flow Conway et al. (2004), which are analogous to the well-known instability of liquid flows in the Taylor-Couette geometry Taylor (1923). The formation of heaps on the top surface of the granular material was also observed in a similar geometry Cabrera and Polanía (2020). In slower flows, where density variations are much smaller, density-driven secondary convection was studied in a cylindrical Couette device, as well as numerically in a split bottom shear cell Krishnaraj and Nott (2016); Dsouza et al. (2017); Dsouza and Nott (2021). Recently, a slow transverse flow was created by a bulldozing mechanism in driving granular material with a conveyor belt Escobar et al. .
@@
Secondary flows are also present in flows of granular suspensions Guazzelli and Pouliquen (2018), which share part of their rheological behavior with dry granular systems Boyer et al. (2011); Trulsson et al. (2012). Secondary flows, possibly associated with migration of particles, are known to be induced by normal stress differences Ramachandran and Leighton (2008), as exemplified by the Weissenberg effect Yamamoto (1958); Oldroyd (1958); Kundu (1973); Elson et al. (1982) with viscoelastic fluids such as polymers. Normal stress differences in granular suspensions have been measured in experiments analyzing the curvature of the free surface of flows in a tilted trough Couturier et al. (2011), or recording stress profiles in a torsional flow between discs Dbouk et al. (2013), as well as in simulations Seto and Giusteri (2018). Interestingly, they get enhanced values for elongated particles such as fibers or dimers Snook et al. (2014); Bounoua et al. (2016); Butler and Snook (2018); Mari (2020). The free-surface curvature of channelized gravity-driven flows of dry grains has also been associated with the presence of normal stress differences McElwaine et al. (2012); Gadal et al. , especially in the context of the formation of levees Félix and Thomas (2004); Deboeuf et al. (2006); Rocha et al. (2019).
@@
The flow of dense granular systems has been largely studied over the last two decades, in particular developing an empirical frictional rheology, known as the $\mu(I)$ rheology GdR MiDi (2004); da Cruz et al. (2005); Jop et al. (2006); Forterre and Pouliquen (2008); Jop (2015). Starting from the idealized case of systems of mono-sized hard spheres, this rheology has been extended in various ways, for example, accounting for grain cohesion Rognon et al. (2006); Mandal et al. (2020); Blatny et al. (2024), grain softness and contact friction Favier de Coulomb et al. (2017), or active (self-propelled) grains Peshkov et al. (2016). An important piece of work has also been done to properly accommodate, or regularize, this rheology in Navier-Stokes solvers Barker et al. (2015, 2017); Goddard and Lee (2018). In case of the presence of strong gradients, a non-local rheology has been developed Pouliquen and Forterre (2009); Bouzid et al. (2013); Henann and Kamrin (2013), for which an additional variable, the fluidity, has to be defined and modeled Bouzid et al. (2015); Zhang and Kamrin (2017); Poon et al. (2023). Finally, the kinetic theory for granular gases, pushed to the dense limit, recovers the features of the granular rheology and its non-local extensions Berzi (2024).
@@
For our present purpose, a key rheological aspect is to account for grain elongation, as e.g. with rods Börzsönyi et al. (2012a); Guo et al. (2013); Berzi et al. (2022), multisphere particles Börzsönyi et al. (2012b); Liu et al. (2024); Rahim et al. (2024), spherocylinders or ellipses Nagy et al. (2017); Trulsson (2018); Nagy et al. (2020). These studies showed that, in a similar way to granular suspensions, these systems display enhanced normal stress differences. In our wish to understand the processes at work in the heaping problem, and to relate secondary transverse vortices to these normal stress differences, the idea is to investigate the flow of grains of arbitrary elongation aspect ratio $\alpha$ in the simpler configuration of the shear cell, where a non-homogeneous forcing is implemented in the flow direction. This forcing will be of arbitrarily small amplitude $\epsilon$ so that we can compute the response of the system at the linear order in this perturbation.
@@
The structure of the paper is the following. In Sec. II we propose a generalized granular rheology in the case of elongated grains and study the case of a plane Couette flow. In Sec. III and IV we analytically compute the linear response of a shear flow to a nonhomogeneous forcing in the flow direction. We introduce the discrete element method (DEM) and numerical setup to simulate such granular flows in Sec. V. In Sec. VI we present and discuss the results of these simulations in comparison to the theoretical expectations.
@@ II.1 Reference isotropic case
Assuming pressure isotropy, it has been proposed by Jop et al. Jop et al. (2006) that the granular rheology, which relates the stress tensor $\sigma_{ij}$ to the velocity field $u_{i}$, takes the form
@@
where $\Pi=\tfrac{1}{3}\sigma_{ii}$ is the pressure and $\dot{\gamma}_{ij}=\dot{\Gamma}_{ij}+\dot{\Gamma}_{ji}$ with $\dot{\Gamma}_{ij}=\partial_{i}u_{j}$ is the shear rate. The norm of the shear rate is here defined as $\dot{\gamma}^{2}=\tfrac{1}{2}\dot{\gamma}_{ij}\dot{\gamma}_{ij}$. The effective friction $\mu$ is a function of the inertial number $I=\dot{\gamma}d/\sqrt{\Pi/\rho}$, where $d$ is the grain size and $\rho$ the grain mass density, and is empirically determined by measurements on experimental or numerical data. Although this rheology also comes with a second constitutive relation describing the behavior of the packing fraction $\phi(I)$, this expression (1) must be consistent with the definition of $\Pi$, incompressibility: $\dot{\gamma}_{ii}=0$. We will eliminate this possible issue by assuming later the quasistatic limit $I\to 0$.
Figure 1: Schematic drawing of the studied system. (a) The simulated cell has a dimension $L_{x}\times L_{y}\times L_{z}$. It is confined by $\sigma_{yy}$ in the vertical direction. Shear is applied in the $x$ direction, plus (b) an additional external force $F_{x}(y,z)$ in the $x$ direction (its magnitude and sign are shown by the surface and the green arrows). This results in the flow pattern shown in (c), four vertices in the velocity field in the $y$-$z$ plane.
@@ II.2 Generalized constitutive relation
We would like to generalize this rheology (1) to the case of elongated grains, for which finite normal stress differences $N_{1}=\sigma_{xx}-\sigma_{yy}$ and $N_{2}=\sigma_{yy}-\sigma_{zz}$ arise. To do so, one must notice that the following symmetries are required
(i) The normal stress differences depend on the direction of the velocity gradient and the flow, so they must change with the transformation $\dot{\bm{\Gamma}}\rightarrow\dot{\bm{\Gamma}}^{T}$. This means they must depend explicitly on $\dot{\Gamma}_{ij}$, as $\dot{\gamma}_{ij}$ is symmetric.
(ii) The normal stress differences are invariant to the reversal of the flow direction ($\dot{\Gamma}_{ij}\rightarrow-\dot{\Gamma}_{ij}$), so they can only depend on even order terms of the shear rate.
(iii) The new term should have a trace of zero, to keep the definition of $\Pi$ correct.
The simplest relation that fulfills these requirements is
@@
or, without index notation, in a coordinate-independent form
@@
with $\dot{\gamma}^{2}=\tfrac{1}{2}\operatorname{Tr}(\dot{\bm{\gamma}}\dot{\bm{\gamma}}^{T})$. The last two terms must have the same factor $\eta$ to ensure that the stress tensor is symmetric. The coefficients $\mu$, $\nu_{1}$, $\nu_{2}$, and $\eta$ depend on the inertial number $I$, but also on microscopic parameters such as the grain aspect ratio $\alpha$ or the grain contact friction $\mu_{p}$ Nagy et al. (2017, 2020); Srivastava et al. (2021). This rheological form is similar to a Rivlin-Ericksen stress closure, as proposed for granular flows in McElwaine et al. (2012); Srivastava et al. (2021); Kim and Kamrin (2023); Gadal et al. .
@@ II.3 Plane Couette flow
The granular rheology is usually studied and calibrated in the geometry of a plane Couette flow. In this case, we consider a cell sheared in the $x$ direction with the gradient in the $y$ (”vertical”) direction, and $z$ being the neutral (”transverse”) direction. This cell has dimensions $L_{x}$, $L_{y}$, and $L_{z}$ along those three axes, and we take the origin of the coordinates at the center of the box. The grains in this cell are subjected to a given pressure $\Pi_{0}$. The flow in the homogeneous case, which will be the base state of the calculation later on, is characterized by a linear velocity profile, associated with a reference shear rate $\dot{\gamma}_{0}$:
@@
In this configuration for which only $\dot{\Gamma}_{yx}\neq 0$, both $\dot{\bm{\Gamma}}_{0}^{2}$ and $(\dot{\bm{\Gamma}}_{0}^{T})^{2}$ are zero, so that the second line of Eq. (3) disappears and the coefficient $\eta$ cannot be measured in such a configuration. To measure it, a configuration with extensional or compressional flows is required. It will then be left adjustable in the following. In contrast, $\nu_{1}$ and $\nu_{2}$ are related to the calibrated first and second normal stress differences as
@@
whose values are taken from Nagy et al. Nagy et al. (2020) (see also Fig. 5 in Appendix A). The traces are related as
@@
As mentioned above, we will restrict to the quasistatic limit for which $I_{0}=\dot{\gamma}_{0}d/\sqrt{\Pi_{0}/\rho}\ll 1$. Under this assumption, the stress tensor in this base state finally reads
@@
where $\pi_{c}=1-\tfrac{1}{3}(\nu_{1c}+\nu_{2c})$, and where the subscript ”c” indicate that $\mu$, $\nu_{1}$ and $\nu_{2}$ are taken constants (independent of the shear rate, but still functions of $\alpha$ and $\mu_{p}$, see Appendix A).
@@ III Conservation equations and perturbation
The equations that govern the dynamics of the system are the conservation of mass and momentum. Assuming incompressibility ($\phi\sim\phi_{c}$ is constant), these equations read
@@
Here, $b_{i}$ is an external body force density. It is an arbitrary perturbation. We take it in the following form:
@@
where $\epsilon$ is its vanishing dimensionless amplitude and $f$ its normalized profile in the $y$-$z$ plane. We take no component along the $y$ and $z$ directions.
Assuming that $\epsilon\ll 1$, the response of velocity and pressure to this perturbation can be written in linear order with respect to the base state as
@@
where $U_{i}$ and $P$ are unknown functions that we wish to relate to $f$. With these expressions, one can compute the shear rate
@@
the symmetric shear rate
@@
and the second-order terms
@@
Even if $\dot{\bm{\Gamma}}^{2}$ and $(\dot{\bm{\Gamma}}^{T})^{2}$ now do not vanish, their traces are still zero to the second order in $\epsilon$, so that
@@
Finally, the stress tensor is
@@
where, using Eq. (3) and the above expressions, we obtain
@@
Figure 2: Comparison of the DEM simulations (red) and calculations based on the generalized granular rheology (blue) for $\alpha{=}2$, $\mu_{p}{=}0.1$, and $L_{z}/L_{y}{=}1$ [presented also in Fig. 4(b)]. (a) The vortex pattern of the $y$-$z$ plane obtained by the coarse-grained averaging of the velocities in simulations, and the expected pattern according to the theory. The width of the arrows represents the magnitude of velocities. (b) Cross sections of the velocity profiles at $y=3.5$ (continuous lines) and $z=7.5$ (dashed lines). In the case of simulations, we present two cross sections obtained by applying different perturbation strengths. Stronger perturbation corresponds to the light red (thicker lines, $A=2\times 10^{-4}$), while weaker perturbation is represented by the dark red (narrower lines, $A=1\times 10^{-4}$).